Ticket Number: HEA-NOC/20101126-1 Ticket Status: CLOSE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20101126 15:29 UTC Problem Start: 20101126 17:11 UTC
Ticket Closed: 20101130 10:19 UTC Problem End: 20101126 17:12 UTC
Site/line: EPA, Co-location
Renumbering HSRP for EPA colo subnet
Brief outage to all hosts in Blanchardstown PoP while HSRP is reconfigured
on HEAnet's routers.
20101126 15:26 OMG Spoke to Frank regarding adjusting their HSRP set up in
Blanchardstown. Changes will be made at 5pm today to bring their set up in
line with other colo customers.
20101126 17:22 OMG Changes made, no problems observed.
20101130 10:19 OMG Closing ticket.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20101126-1
HEAnet Limited. Registered in Ireland, No. 275301.